OpenStudy (anonymous):

Solve the differential equation y prime equals the product of 4 times x and the square root of the quantity 1 minus y squared. b. Explain why the initial value problem y prime equals the product of 4 times x and the square root of the quantity 1 minus y squared with y(0) = 4 does not have a solution.

OpenStudy (irishboy123):

\(y' = 4 \sqrt{1 - y^2}\) so \(\int \dfrac{dy}{4 \sqrt{1 - y^2}} = \int dx \) do you mean that, or summat altogether different?

OpenStudy (anonymous):

yes that

OpenStudy (irishboy123):

from there, it's easy